OpenStudy (anonymous):

Find the slope of the line whose equation is 4x – 6y = 12

OpenStudy (anonymous):

@DarkBlueChocobo

OpenStudy (anonymous):

@Kit_Kat_Nat_<3

OpenStudy (anonymous):

Ok, this is in slope intercept form so do you know what the slope would be? y=mx+b m=slope

OpenStudy (anonymous):

change the equation to the form y=mx+b

OpenStudy (darkbluechocobo):

4/1 right?

OpenStudy (darkbluechocobo):

well dang it :p

OpenStudy (darkbluechocobo):

i havent done this since geometry ._.

OpenStudy (anonymous):

Lol @DarkBlueChocobo

OpenStudy (anonymous):

and your m would be the gradient which is also the slope

OpenStudy (darkbluechocobo):

Lelelel and now im in pre-cal the life :S

OpenStudy (anonymous):

WAIT IM SO CONFUSED

OpenStudy (darkbluechocobo):

so basically

OpenStudy (darkbluechocobo):

rearrange the equation

OpenStudy (darkbluechocobo):

so that it looks like mx+b=y

OpenStudy (anonymous):

OK

OpenStudy (darkbluechocobo):

soh in this case If I am correct. mx=4 b=12 y=6

OpenStudy (cwrw238):

rearrange: 6y = 4x - 12 now divide by 2 3y = 2x - 6 now divide by 3 and you have it

OpenStudy (anonymous):

so 4 + 12 = 6?

OpenStudy (darkbluechocobo):

well missing the x's and y's

OpenStudy (darkbluechocobo):

other than that so far so good

OpenStudy (darkbluechocobo):

@cwrw238 dis guy knows whats up :p

OpenStudy (anonymous):

how to u divide 2 by 3

OpenStudy (darkbluechocobo):

it will come out as a decimal

undeadknight26 (undeadknight26):

Use a calculator ._.

OpenStudy (darkbluechocobo):

or you can just put 2/3x

OpenStudy (darkbluechocobo):

works that way too

OpenStudy (cwrw238):

or you can write it as a fraction

OpenStudy (darkbluechocobo):

^^^

OpenStudy (anonymous):

so its 2/3?

OpenStudy (cwrw238):

yep

OpenStudy (darkbluechocobo):

mhm ding ding you be correctarino

OpenStudy (anonymous):

yayyyyyyyy

OpenStudy (darkbluechocobo):

wopwowpwopwp

OpenStudy (anonymous):

i have another math paper with like 5 questions could u help with that too?

OpenStudy (darkbluechocobo):

kk ive got time i got all my work basically done for the week

OpenStudy (anonymous):

ok ill open a new one